You wish to use a spring that has a constant of 25N/m to launch a 25.0g object. If the spring is initially compressed 5.0cm before launching the projectile, what is the maximum height the object will attain.


1
Expert's answer
2022-01-13T13:25:10-0500

by energy conservation:

"\frac{kx^2}{2}=mgh"


then, the maximum height the object will attain:

"h=\frac{kx^2}{2mg}=\frac{25\cdot 0.05^2}{2\cdot 25\cdot10^{-3}\cdot 9.8}=0.13" m